# LLM4Cell: A Survey of Large Language and Agentic Models for Single-Cell Biology

### v1

@@ 1 Introduction
Large language models (LLMs) are transforming biomedical discovery by linking molecular patterns with knowledge encoded in text. In single-cell biology, where each experiment measures thousands of genes across millions of cells, LLMs promise to unify gene expression, chromatin accessibility, and spatial organization under a shared, interpretable framework. Early foundation models such as scGPT  (Cui et al., 2024), Geneformer  (Theodoris et al., 2023), and scFoundation  (Hao et al., 2024) learn transferable representations directly from single-cell data, while emerging systems like CellLM  (Zhao et al., 2023), scAgent  (Mao et al., 2025), and CellVerse  (Zhang et al., 2025b) extend these capabilities toward reasoning, dialogue, and autonomous analysis. Together, they signal a shift from statistical inference to language-driven, interpretable, and generalizable cell intelligence.
Despite this progress, the field remains fragmented. Models differ widely in data modality, supervision type, and evaluation standard—ranging from scRNA-seq to ATAC, multi-omic, and spatial measurements—hindering cross-model comparison and reproducibility. Benchmarking is inconsistent, datasets are unevenly distributed across modalities, and agentic frameworks lack standardized ways to measure reasoning correctness or biological grounding. Without a systematic synthesis, it is difficult to assess what current single-cell LLMs truly understand or where they fall short.
To address these challenges, we introduce LLM4Cell, a comprehensive survey of large language and agentic models for single-cell biology. We analyze 58 representative methods spanning five methodological families—Foundation, Text-Bridge, Spatial/Multimodal, Epigenomic, and Agentic—and organize them across eight major tasks, from annotation and trajectory inference to perturbation and drug-response modeling. Complementing this, we compile over 40 publicly available datasets across RNA, ATAC, multi-omic, spatial, perturbation, and plant domains, and evaluate each model using a ten-dimension rubric covering grounding, fairness, scalability, and interpretability.
@@
LLM4Cell provides: (1) a modality-balanced registry of $\sim$40 benchmark datasets for large-model training and evaluation; (2) a unified taxonomy of 58 foundation and agentic models across eight analytical tasks; (3) a ten-dimension domain rubric capturing biological grounding, generalization, and ethics; and (4) a critical discussion of open challenges in cross-modal alignment, reasoning, and trustworthy AI for cell biology.
@@ 2 Related Work
However, existing reviews remain limited in scope. They focus on architectures or prompting but give little attention to the datasets underpinning model training and evaluation—especially those covering spatial, perturbation, or multimodal modalities. Standardized benchmarks, domain-aware metrics, and reproducible splits are seldom addressed, and plant single-cell datasets—key for cross-kingdom generalization—are almost entirely absent. Moreover, agentic reasoning and tool-augmented systems are typically treated as isolated examples rather than a coherent methodological class, leaving the relationship among data modalities, modeling paradigms, and biological domains fragmented.
@@ 3 Datasets
Progress in large language models for single-cell biology relies on the rapid growth of high-quality datasets across transcriptomic, epigenomic, multimodal, and spatial domains. Our survey compiles over forty public resources spanning five major modalities—RNA, ATAC, multi-omic, spatial, and perturbation/drug-response—plus emerging plant single-cell atlases. These datasets underpin model pretraining, evaluation, and cross-modal reasoning for annotation, integration, trajectory inference, perturbation modeling, and spatial mapping.
@@ 4 Model Taxonomy
Large language models are rapidly reshaping single-cell biology, producing diverse architectures, pretraining schemes, and reasoning frameworks. We categorize 58 representative methods into five methodological families—Foundation Models, Text-Bridge LLMs, Spatial and Multimodal Models, Epigenomic Models, and Agentic Frameworks—based on their core design and data modality. These span the progression from gene-level embeddings to multimodal and autonomous systems capable of biological reasoning.

### v2

@@ 1 Introduction
Large language models (LLMs) are transforming biomedical discovery by linking molecular patterns with knowledge encoded in text.  (Zhang et al., 2025a; Aljohani et al., 2025; Bi et al., 2024). Prior work on computational workflows and biological prediction tasks demonstrates how learned representations capture complex molecular relationships and enhance practical performance.  (Lam et al., 2024; Bhattacharya et al., 2024; Dip et al., 2024; Nam et al., 2024; Shuvo et al., 2024) In single-cell biology, where each experiment measures thousands of genes across millions of cells, LLMs promise to unify gene expression, chromatin accessibility, and spatial organization under a shared, interpretable framework. Early foundation models such as scGPT  (Cui et al., 2024), Geneformer  (Theodoris et al., 2023), and scFoundation  (Hao et al., 2024) learn transferable representations directly from single-cell data, while emerging systems like CellLM  (Zhao et al., 2023), scAgent  (Mao et al., 2025), and CellVerse  (Zhang et al., 2025b) extend these capabilities toward reasoning, dialogue, and autonomous analysis. Together, they signal a shift from statistical inference to language-driven, interpretable, and generalizable cell intelligence.
Despite this progress, the field remains fragmented. Models differ widely in data modality, supervision type, and evaluation standard ranging from scRNA-seq to ATAC, multi-omic, and spatial measurements hindering cross-model comparison and reproducibility. Benchmarking is inconsistent, datasets are unevenly distributed across modalities, and agentic frameworks lack standardized ways to measure reasoning correctness or biological grounding. Without a systematic synthesis, it is difficult to assess what current single-cell LLMs truly understand or where they fall short.
To address these challenges, we introduce LLM4Cell, a comprehensive survey of large language and agentic models for single-cell biology. We analyze 58 representative methods spanning five methodological familiesFoundation, Text-Bridge, Spatial/Multimodal, Epigenomic, and Agentic and organize them across eight major tasks, from annotation and trajectory inference to perturbation and drug-response modeling. Complementing this, we compile over 40 publicly available datasets across RNA, ATAC, multi-omic, spatial, perturbation, and plant domains, and evaluate each model using a ten-dimension rubric covering grounding, fairness, scalability, and interpretability.
@@
LLM4Cell provides: (1) a modality-balanced registry of around 40 benchmark datasets for large-model training and evaluation; (2) a unified taxonomy of 58 foundation and agentic models across eight analytical tasks; (3) a ten-dimension domain rubric capturing biological grounding, generalization, and ethics; and (4) a critical discussion of open challenges in cross-modal alignment, reasoning, and trustworthy AI for cell biology.
@@ 2 Related Work
However, existing reviews remain limited in scope. They focus on architectures or prompting but give little attention to the datasets underpinning model training and evaluation especially those covering spatial, perturbation, or multimodal modalities. Standardized benchmarks, domain-aware metrics, and reproducible splits are seldom addressed, and plant single-cell datasets key for cross-kingdom generalization are almost entirely absent. Moreover, agentic reasoning and tool-augmented systems are typically treated as isolated examples rather than a coherent methodological class, leaving the relationship among data modalities, modeling paradigms, and biological domains fragmented.
@@ 3 Datasets
Progress in large language models for single-cell biology relies on the rapid growth of high-quality datasets across transcriptomic, epigenomic, multimodal, and spatial domains. Our survey compiles over forty public resources spanning five major modalities RNA, ATAC, multi-omic, spatial, and perturbation/drug-response plus emerging plant single-cell atlases. These datasets underpin model pretraining, evaluation, and cross-modal reasoning for annotation, integration, trajectory inference, perturbation modeling, and spatial mapping.
@@ 4 Model Taxonomy
Large language models are rapidly reshaping single-cell biology, producing diverse architectures, pretraining schemes, and reasoning frameworks. We categorize 58 representative methods into five methodological families Foundation Models, Text-Bridge LLMs, Spatial and Multimodal Models, Epigenomic Models, and Agentic Frameworks based on their core design and data modality. These span the progression from gene-level embeddings to multimodal and autonomous systems capable of biological reasoning.